OpenStudy (anonymous):

Evaluate the expression without using a calculator. 10log(x^3)

OpenStudy (anonymous):

30 log x

OpenStudy (anonymous):

\[10 \log(x^3) = 3\times10\log(x) = 30\log(x)\]

OpenStudy (anonymous):

Thanks, both of you